MONT 107Q – Thinking About Mathematics Study and Discussion Question on Ovid, The Metamorphoses March 13, 15, 17, 2017

Book I

1. How does the poet Ovid’s “book of changes” fit with this year’s CHQ theme? What borders are involved and how do the stories here involve borders?

2. How does the creation story at the beginning of Book I (pages 3 - 6) compare with the creation story in the Book of Genesis from the Bible? What are the similarities? What are the differences?

3. The description of the four ages (pages 6 - 8) describes human history as a continual descent from a better past into a worse present. Our own Western culture’s selfnarrative, over the past several centuries at least, has generally been one going the opposite way – a story of continual progress and improvement. Which of those trajectories seem to match the way human societies are headed in 2017? What does how we tell our story say about us?

4. The events leading up to the story of Deucalion and Pyrrha (pages 9 - 19) also have Biblical parallels. What story does this section of Ovid’s poem relate to? Something to think about: Could myths like this from many cultures preserve distant memories of real events?

5. The Pythian Games (p. 20) were actual events in ancient Greece–one of the series of four recurring pan-Hellenic (all-Greek) games. Where were they held? What was the other cultural focus at this location? (Think back to Herodotus last fall!)

6. Why does Ovid turn to the myth of Apollo and Daphne (pages 20 - 25) immediately after the slaying of Python? What is the connection? What does this suggest about the structure of the poem as a whole? Which god actually does Daphne’s metamorphosis? What is the symbolism involved in this story? How would you interpret it?

7. The story of Apollo and Daphne has also been a powerful stimulus for artists in various media. Look up the statue “Apollo and Daphne” by the Italian sculptor (and painter, and architect) Gian Lorenzo Bernini (1598 - 1680). What is being depicted there?

8. In the next story, Jove and Io, (pages 26 - 30, 31 - 33), Io undergoes two transformations – how does that work and who does each of them? Note that the story of Pan and Syrinx (pages 30 - 31) is effectively folded into the Jove-Io story! What is the connection? Who is Argus? What becomes of him? Do we see a different sexual dynamic here than in the Apollo and Daphne story?

Book II

1. The final tale of Book I (the story of Phaethon) continues at the start of Book II – (pages 33 - 49). Is the point of this story simply the warning “don’t bite off more than you can chew,” so to speak? Or does Ovid also make us question that moral? Think about the epitaph of Phaethon (page 49). Also think about the follow-up stories of the Heliades (page 50) and Cycnus (page 51). By mourning Phaethon (excessively, maybe?) what becomes of them? How does Phaethon’s father react to the whole regrettable incident? (page 52)

2. What are the two transformations that Callisto undergoes (pages 52 - 57), one at the hand of Juno, and the second at the hand of Jove? What does it mean that the Bear “never taints the stream of pure Ocean?”

3. The middle section of Book II is a series of nested, interlocking stories (The Raven and Crow, Ocyrhoe, Mercury and Battus, the Envy of Aglauros). Are there common themes that link them?

4. The story of Jupiter and Europa (pages 71 - 73) is similar to several others in that Jove (Jupiter) seduces a human woman. But are there differences as well? Has the earlier story of Callisto perhaps changed him? (Hint: the story continues at the start of Book III!)

5. One might say that Book II as a whole deals with the relationship between humans and nature (or humans and the gods, in the understanding of the Greeks and Romans). How do each of the stories here relate to that theme?

Book III

1. How effective is the transition between Book II and Book III? (Recall the statement by Ovid at the beginning that he wants his song to be “seamless.”)

2. What are the main themes of the stories in Book III? Is the relationship between men and gods different here than in Book II?

3. The Story of Actaeon (pages 81 - 86): What is Actaeon’s attitude toward “the hunt”? What are he and his friends like as hunters? Do you think Actaeon’s punishment is fair?

4. What happens to Semele (pages 86 - 89) ? Who is her child and how is it that he is twice born?

5. What is the humorous controversy between Jove and Juno about the sexual act? Why do they ask Tiresias to decide? (pages 89 - 90)

6. What is the purpose of the story of Narcissus? How does the motif of seeing link the stories (for example Tiresias, pages 89 - 90) and Narcissus and Echo (pages 90 - 97)?

7. Cadmus’ grandson Pentheus scorns the seer Tiresias and he also scorns the god Bacchus. What does Pentheus fear? (pages 97 - 106)

8. Is there a difference between Actaeon’s and Pentheus’ transgressions. Is one a sin? Is one an error? Is there a difference?

9. Are there as many instances of physical metamorphosis in this book? Are most of the changes of other sorts?

Book XV

1. About half of this book is taken up by a long speech by the philosopher Pythagoras (pages 515 - 532). What is the function of this and how does it related to the rest of the Metamorphoses? What are the major aspects of Pythagoras’ teachings that Ovid focuses on? Is Ovid trying to say that a famous philosopher agrees with him, so he (Ovid) must be right?

2. What is Ovid trying to do with the very last section (pages 548 - 549)? 3